# Hotfixing Large Language Models for Code: How Far Can Parameter-Efficient Fine-Tuning Go?

### v2

@@ 3.2. Data Collection
Generally, the training data for hotfixing should include three types of code: (1) the desired code that we want the model to generate, (2) the undesired code that we want the model to avoid, and (3) the code that is not relevant to the two types of code. Such information can be retrieved from the diff between two versions of one code snippet. One version is the code snippet that we want the model not to generate (e.g., buggy code). The other version is the code snippet that we want the model to generate (e.g., the corresponding fixed code). Computing the diff can help find what code should be changed and what code should remain the same.
@@
We reuse the SStuBs dataset provided by Jesse et al. (Jesse et al., 2023), which is also known as ManySStuBs4J (Karampatsis and Sutton, 2020), coming from the 2021 MSR mining challenge dataset. An example from the dataset is shown in Listing 1. The code highlighted in red is the buggy code. The code highlighted in green is the corresponding fix. This dataset consists of two parts: a small dataset and a large dataset. The large dataset, which includes 63,923 single-statement bug-fix changes, is chosen in the experiment. After removing duplicates for bugs that share the exact same prefix (i.e., the content before bug locations), bug, and fix, Jesse et al. conduct their evaluation on 16,899 examples. We reuse their filtered dataset in our experiment. We split the dataset into training, validation, and test sets with a ratio of 8:1:1. The training set is used to hotfix the models. The test set is used to evaluate number of generated fixed code and buggy code before and after hotfixing.
@@
In Jesse et al.’s experiment, the content before the buggy code is sent to a code generator as the prompt. Given a prompt, the model output has three possible cases: (1) containing the buggy code, (2) containing the fixed code, or (3) not matching any of the two. By sending many prompts to a model, Jesse et al. count the number of cases (1) and (2) and compute their ratio to show that models tend to generate buggy code. As these models are non-deterministic, the same prompt may lead to different outputs. We let the model generate ten outputs for each prompt and count the number of cases (1) and (2).
3.3. Learning Objectives
Recall how LLM4Code generate outputs: given a sequence of tokens as input, the model generates a distribution over the next token. Given a training example input $x$ and its targeted output $y$, the model computes the probability of outputting $y$ given $x$, denoted by $P(y|x)$. The model learns this training example by maximizing this probability. In practice, this processes is conducted by minimizing the negative log-likelihood (NLL) loss function, which is defined as:
(2) $L_{vanilla}=-\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}|x_{i})$
where $N$ is the number of training samples, $x_{i}$ is the input for the $i$-th sample, $y_{i}$ is the target output for the $i$-th sample. This loss function is commonly used to tune LLM4Code (Weyssow et al., 2024). By minimizing this loss function, the model adjusts its parameters to increase the probabilities of the target words, improving its ability to predict the next word in similar sentences. We call it the Vanilla loss function.
However, in the context of hotfixing LLM4Code, directly training models on the entire updated examples may not be the appropriate option. The reason is highlighted in the work by He and Vechev (He and Vechev, 2023). In hotfixing, not each token in the entire updated examples is equally important, and the model should pay more attention to the added fixed code. Besides, the model should also learn not to generate the original buggy code and the model should cause minimal impact on the outputs that are not relevant to the bug-fixing updates. We explain the three objectives and the corresponding loss function design as follows.
3.3.1. Learning to Generate the Desired Code
In Listing 1, the line that is highlighted in green is the output that we want the model to generate. To specifically control that such learning only happens on the desired code (i.e., the added fixed code), we compute the loss on the desired part only. We apply Equation (2) to compute the loss:
(3) $L_{guided}=-\frac{1}{N}\sum w_{t}^{+}\log P(x_{t}|x_{<t})$
In the above equation, $w_{t}^{+}$ is the weight to compute the final loss on token $x_{t}$, where $w_{t}=1$ if $x_{t}$ is the desired code and $w_{t}^{+}=0$ otherwise. This loss setting is called the Guided loss.
3.3.2. Learning to Avoid Undesired Code
In Listing 1, the part that is highlighted in red (e.g., the buggy code) is the output that we want the model to avoid. Similarly, the log-likelihood loss function on the undesired code can be computed as follows:
(4) $L_{unlearn}=-\frac{1}{N}\sum w_{t}^{-}\log P(x_{t}|x_{<t})$
In the above equation, $w_{t}^{-}$ is the weight in computing the loss on token $x_{t}$, where $w_{t}^{-}=1$ if $x_{t}$ is the desired code and $w_{t}^{-}=0$ otherwise. Unlike the desired code, we want to increase the model loss on the undesired code to avoid them from being generated. However, simply maximizing the loss on the undesired code is ineffective. Our experiment shows that maximizing this loss will make the model lose its ability in generating meaningful code (the hotfixed models generate zero buggy code and zero fixed code), which is not desired in hotfixing. Instead of directly maximizing this loss, He and Vechev (He and Vechev, 2023) propose to adopt the loss function $\frac{L_{guided}}{L_{guided}+L_{unlearn}}$, which guides the optimizer to find a balance between minimizing the loss on desired code $L_{guided}$ and maximizing the loss on undesirable code $L_{unlearn}$. Ideally, to minimize this loss, the optimizer works to reduce $L_{guided}$ which encourages fixed code generation and to increase $L_{unlearn}$ which discourages the generation of buggy code. To make the model learn the desired code and avoid the undesired code, we define the Dual loss as follows:
(5) $L_{dual}=(L_{guided}+\frac{L_{guided}}{L_{guided}+L_{unlearn}})/2$
3.3.3. Learning to Retrain Knowledge of Other Code
The third objective is to let the model preserve the knowledge of the code that is not relevant to either the desired or undesired code. Although the loss functions designed above do not touch the other code, the model may forget the knowledge of this part during training, which may introduce other undesired behaviors (e.g., harming the correctness of generated code). As an analogy, a commit fixing a bug may introduce another. We want to avoid such a situation for hotfix.
The fact that tuning language models on a new dataset can lead to performance degradation on other aspects has been observed in previous studies. For example, Liu et al. (Liu et al., 2016) point out that optimizing models on a new dataset can improve the average performance (e.g., BLEU score) but forgo coherence and fluency. This degeneration is often diagnosed as an effect of deviating too much from the original pre-trained model during optimization. Consequently, we want the model to behave similarly to the original model on the code that is not relevant to the undesired behavior. This is similar to the target of knowledge distillation (Tang et al., 2024; Shi et al., 2024), which aims to train a student model to mimic the behavior of the teacher model. In knowledge distillation, Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) is widely used to measure the behavioral difference between the two models and the student model learns to minimize the difference (Shi et al., 2023; Kim and Rush, 2016; Kim et al., 2024). He and Vechev also use it to mitigate the potential functional correctness decrease. The KL divergence loss is defined as follows:
(6) $L_{KL}=\sum D_{KL}(P(x_{t}|x_{<t})||P_{0}(x_{t}|x_{<t}))$
In the above equation, $P_{0}$ is the probability generated by the original model. In Equation (6), the notation $D_{KL}$ represents KL divergence between the original model’s token distribution and the hotfixed model’s token distribution. KL divergence is a measure of how one probability distribution diverges from a second, expected probability distribution. A small KL divergence indicates that the two distributions are similar. In our experiment, we also add $L_{KL}$ to each of the loss functions designed above, denoted by Vanilla+KL, Guided+KL, and Dual+KL. For example, the loss function of Guided+KL is $(L_{guided}+L_{KL})/2$ and the loss function of Dual+KL is $(L_{guided}+\frac{L_{guided}}{L_{guided}+L_{unlearn}}+L_{KL})/3$
3.4. Fine-tuning LLM4Code
Another important requirement for hotfixing LLM4Code is to make the process efficient and easy to deploy. Considering the efficiency requirement, we employ various parameter-efficient fine-tuning (PEFT) techniques to update the models, which can reduce the number of trainable parameters and thus reduce the computational cost. Following a recent study on evaluating PEFT on LLM4Code (Weyssow et al., 2024), we employ four PEFT techniques: LoRA (Hu et al., 2022b), IA3 (Liu et al., 2022), Prefix tuning (Li and Liang, 2021), and QLoRA (Dettmers et al., 2023). For QLoRA, we quantize the model to 8-bit floating-point and 4-bit floating-point data types. In the following part of the paper, we use the acronym FT, LoRA, IA3, Prefix, QLoRA-8bit, and QLoRA-4bit to represent the traditional fine-tuning, LoRA, IA3, Prefix tuning, QLoRA with 8-bit quantization, and QLoRA with 4-bit quantization, respectively. Note that the hotfix is in different forms depending on the PEFT methods used. For example, the hotfix using prefix-tuning is a set of virtual tokens added to the input tokens of the LLM.
4. Experiment Settings
This section explains the investigated models, baselines, potential side effects brought by hotfixing, and model training details.
4.1. Investigated Models
@@
Considering the popularity and high performance of decoder-only models in code generators (Hou et al., 2023), we focus on decoder-only models in our evaluation. Following the study that exposes the undesired behaviors of code generators (Jesse et al., 2023), we choose CodeGen (Nijkamp et al., 2023) models as our experiment subject, which are open-source models and have been widely used when evaluating code generation models. The CodeGen architecture follows a standard transformer decoder with left-to-right causal masking. Nijkamp et al. (Nijkamp et al., 2023) train multiple CodeGen models using different configurations and datasets. Although CodeGen-mono achieves the best performance on the HumanEval benchmark, it only supports the generation of Python code. We choose CodeGen-multi, which can support multiple programming languages, as our experiment subject. We use three CodeGen-multi models of different sizes in terms of the number of parameters: 350M and 2B. [^4] In this paper, we use CodeGen-350M and CodeGen-2B to refer to the CodeGen-350M-multi and CodeGen-2B-multi. Considering that we need to train multiple models with different fine-tuning methods, we exclude the larger version with 6 and 16 billion parameters due to the limit of our computational resources.
4.2. Baseline
To encourage the models to generate fixed code, Jesse et al. (Jesse et al., 2023) add a natural language comment before the location of bugs to pre-condition the model. They use a CodeTrans model (Elnaggar et al., 2021) fine-tuned on the DeepCom dataset (Hu et al., 2018) to generate comments for each buggy statement and its corresponding fix, which we call BugComment and FixComment, respectively. The experiment results (Jesse et al., 2023) show that FixComment can condition the model to generate more fixed code. For example, after adding FixComment, CodeGen-350M (Nijkamp et al., 2023) generates 7.58% less buggy code and 13.12% more fixed code. We use this strategy as a baseline to compare with hotfixing. Note that to generate the comment, the baseline assumes the knowledge of the fix to be completed each time it mitigates the undesired behavior, which is not always available in practice. In contrast, hotfixing does not require the such knowledge of the fix when it is used at the model inference time.
@@ 4.3.2. Perplexity
$N$ is the total number of words in the sequence. $P(w_{i}\mid w_{1},\ldots,w_{i-1})$ is the probability assigned by the model to the word $w_{i}$ given the preceding words $w_{1},\ldots,w_{i-1}$. A lower perplexity score indicates that the knowledge of the model is more consistent with the data used to compute the perplexity. We aim to evaluate how hotfixing may affect the model’s knowledge about the original training data. However, the training data of both CodeGen-350M and CodeGen-2B are not publicly available. Given the insights from previous studies that larger models can memorize much training data (Yang et al., 2024b; Carlini et al., 2021), we sample 10,000 outputs from the CodeGen-16B model, which are expected to be similar to the training data of CodeGen-350M and CodeGen-2B. Thus, these sampled outputs are appropriate to evaluate how model’s knowledge about the original training data is affected by hotfixing. We measure the average perplexity of CodeGen-350M, CodeGen-2B, and their hotfixed versions on the 10,000 outputs sampled from CodeGen-16B.
Table 1. This table shows how PEFT methods affect hotfixing. We set the loss function of each model to be the same. The integer number in each cell means the number of generated buggy/fixed code. The percentage in the parentheses shows the change compared to the original models’ results. $\downarrow$’ and $\uparrow$’ indicate the decrease and increase, respectively. The cell of best results are highlighted in green.
@@
4.4. Experiment Settings
We download the CodeGen models from the HuggingFace and run them on two NVIDIA GeForce A6000 GPUs with 48 GB of memory. We obtain the dataset from Jesse et al.’s replication package (Jesse et al., 2023). We also conduct a case study on mitigating privacy leakage in LLM4Code. The required dataset bigcode/bigcode-pii-dataset is downloaded from HuggingFace. [^5] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset We obtain the personally identifiable information (PII) detector bigcode/starpii from HuggingFace. [^6] https://huggingface.co/bigcode/starpii It is noted that neither bigcode-pii-dataset nor starpii is publicly accessible, which requires approval from the dataset owner to access. We extend the open-source repository provided by Weyssow et al. (Weyssow et al., 2024) to implement different fine-tuning methods with our proposed loss functions. Following their settings, we set the learning rate to 3e-4. For prefix-tuning, the prefix is set to be 20 trainable continuous tokens. The Adafactor optmizer (Shazeer and Stern, 2018) with 16-bit floating-point precision is used for training.
5. Results
This section presents three research questions and the corresponding results regarding hotfixing LLM4Code.
RQ1. How do different PEFT methods affect hotfixing results?
RQ2. How do learn objectives affect hotfixing results?
RQ3. How does hotfixing cause side effects to LLM4Code and how to mitigate the undesired side effects?
The first question tries different PEFT strategies and keep the same learning objectives. The second question uses the same fine-tuning method (LoRA) and try different learning targets to understand how they affect the hotfixing results. We also examine how hotfixing causes side effects and how to mitigate the effects.
5.1. RQ1. How do different PEFT methods affect hotfixing results?
@@ 5. Results
In this research question, we hope to understand what PEFT methods can perform better in hotfixing LLM4Code. Here the performance has two aspects: (1) effectiveness: to what extent the hotfix can make the model generate less buggy but more fixed code; and (2) efficiency: how much computational resources are required to hotfix the model.
@@
Table 1 shows the results of this research question. First, the third column shows the number of buggy code and fixed code generated by the original CodeGen-350M and CodeGen-2B models. We observe that both models tend to generate more buggy code than fixed code, confirming the findings in Jesse et al.’s study (Jesse et al., 2023). We also observe that the larger model generates more buggy code and fixed code than the smaller model, indicating that larger models have stronger capacity to memorize the training data (Yang et al., 2024b). Then, we use the mitigation method proposed in (Jesse et al., 2023) as a baseline and find that adding comments to the model prompt can mitigate the issue to some extent but the performance is limited. Specifically, this method only reduces 7.58% and 10.79% of buggy code, and increases 13.12% and 6.38% of fixed code for CodeGen-350M and CodeGen-2B, respectively.
@@
Table 2. The two tables shows how loss functions affect hotfixing. We use LoRA to train each model. The percentage number under ‘Vanilla’, ‘Guided’, and ‘Dual’ columns are changes compared to the original model. The percentage number under ‘Vanilla+KL’, ‘Guided+KL’, and ‘Guided+KL’ columns are the changes compared to the ‘Vanilla’, ‘Guided’, and ‘Dual’ columns. $\downarrow$’ and $\uparrow$’ indicate the decrease and increase.
Type Models Vanilla Guided Dual
@@
Vanilla+KL Guided+KL Dual+KL Changes by KL
@@
We then evaluate the performance of hotfixing under multiple settings. In this RQ, we use the combined loss function to fine-tune the models, so that the model can learn to avoid generating undesired code, generate the desired code, and remain the same for other code. We find that use this loss function can reduce 44.32% (for CodeGen-350M) and 50.47% (for CodeGen-2B) of buggy code, and increase 108.43% (for CodeGen-350M) and 103.38% (for CodeGen-2B) of fixed code, demonstrating the effectiveness of hotfixing. Then, we compare the difference between different PEFT methods. In Table 1, we label a cell as green if a PEFT method achieves the best performance. We can observe that LoRA has 3 green cells and the other one is QLoRA-4bit. It indicates the that LoRA-based methods are more effective in hotfixing LLM4Code. In contrast, IA3 and Prefix-tuning are less effective. For example, prefix-tuning only leads to 5.40% more fixed code generation for CodeGen-2B, which is much less than the 103.38% increase by LoRA.
We also analyze the efficiency of different PEFT methods. We set the batch size to 1 for each method and record the average time required to conduct one epoch of learning on the same A6000 GPU. We repeat the experiment 5 times and report the average time. On CodeGen-350M, LoRA and Prefix-tuning shows better time efficiency, which takes 5.6 and 5.4 minutes per epoch. IA3 takes 6.6 minutes, faster than QLoRA-4bit and QLoRA-8bit that take 7 and 9.6 minutes per epoch, respectively.
Answers to RQ1: LoRA-based methods demonstrate better effectiveness and efficiency in hotfixing LLM4Code than IA3 and prefix-tuning. In our best setting, hotfixing can reduce 44.32% and 50.47% of buggy code and increase 108.43% and 103.38% of fixed code for CodeGen-350M and CodeGen-2B.
5.2. RQ2. How do learning objectives affect hotfixing results?
Recall the objectives when hotfixing LLM4Code: (1) avoid generating undesired code, (2) generate the desired code, and (3) remain the same for other code. In RQ1, we consider all the three objectives and train models with different PEFT methods, identifying LoRA as the best-performing training method in hotfixing. In RQ2, we use LoRA to train models and analyze how each learning objective affects hotfixing results.
In Section 3.3, we define three settings of loss functions, named Vanilla, Guided, and Dual loss settings. In the Vanilla setting, we minimize the loss on the entire example after fix, i.e., the commonly used setting in fine-tuning LLM4Code (Weyssow et al., 2024), defined in Equation (2). In the Guided setting, we guide to model to minimize the loss only on the added fixed code, defined in Equation (3). In the Dual setting, we additionally penalize the loss on the deleted code on top of the Guided setting, defined in Equation (5). The results of each loss setting are shown in the left side of Table 2. It can be observed that all the three settings can mitigate the undesired behaviors of LLM4Code. The Vanilla setting reduces 11.38% and 8.69% of buggy code and increases 64.89% and 48.02% of fixed code for CodeGen-350M and CodeGen-2B, respectively. By guiding the model to specifically learn on the added fixed code, the Guided setting can reduce 37.81% and 41.88% of buggy code and increase 93.11% and 100.90% of fixed code for the two models. Additionally penalizing the model on the deleted code in the Dual setting can further improve the performance. Taking CodeGen-350M as an example, the Dual setting can reduce 41.18% of buggy code and increase 107.05% of fixed code, outperforming the Guided setting that reduces 37.81% of buggy code and increases 93.11% of fixed code. The results suggest that guiding the model to learn the difference between two versions of dataset is more effective in hotfixing.
@@
We also investigate the impact of adding KL divergence to the loss function. The goal of adding KL divergence is to reduce the potential side effects caused by hotfixing (detailed analysis in RQ3). However, this may also affect the hotfixing results and we show that such impacts on hotfixing results are limited. The right half of Table 2 shows the results when adding KL divergence to each loss function, in columns with the suffix ‘+KL.’ The percentage number in this table is the change compared to the results without KL divergence. We find that adding KL loss does not have a significant impact on the hotfixing results. We report the average changes caused by KL loss in the last column of Table 2. On average, adding KL loss only increases 3.41% of buggy code and reduces 1.51% of fixed code for CodeGen-350M, and increases 2.74% of buggy code and reduces 1.57% of fixed code for CodeGen-2B, compared to the results without KL loss. However, as we will show in the next research question, adding KL loss can help mitigate the undesired side effects caused by hotfixing LLM4Code.
@@
Answers to RQ2: Our Dual loss setting can achieve the best performance, reducing 41.18% and 50.47% of buggy code and increasing 107.05% and 103.38% of fixed code for CodeGen-350M and CodeGen-2B, respectively. Adding KL divergence to the loss function does not have a negative significant impact on results.
@@
5.3. RQ3: How does hotfixing affect functional correctness and how to mitigate the undesired side effects?

### v4

@@ 3.4. Data Collection
Software changes are in the form of diff, which contains both the added code (e.g., fixed code) and the removed code (e.g., buggy code). By excluding the added and removed code, we can obtain the code that is not updated during the bug fixing. We reuse the SStuBs dataset provided by Jesse et al. (Jesse et al., 2023), also known as ManySStuBs4J (Karampatsis and Sutton, 2020), originating from the 2021 MSR Mining Challenge dataset. An example from the dataset is shown in Listing 1. The code highlighted in red is the buggy code that is removed. The code highlighted in green is the corresponding fix that is added. The remaining part is untouched in the bug-fixing process.
@@
The ManySStuBs4J dataset (Karampatsis and Sutton, 2020) consists of two parts: a small dataset and a large dataset. The large dataset, which includes 63,923 single-statement bug-fix changes, is chosen for the experiment. After removing duplicates that share the exact same prefix (i.e., the content before bug locations), bug, and fix, Jesse et al. conduct their evaluation on 16,899 examples. We reuse their filtered dataset in our experiment. We split the dataset into training, validation, and test sets with a ratio of 8:1:1. The training set is used to hotfix the models. The test set is used to evaluate the number of generated fixed code and buggy code before and after hotfixing.
@@
In Jesse et al.’s experiment, the content before the buggy code is sent to a code generator as the prompt. As the ManySStuBs4J dataset is a collection of single-statement bug-fix changes, we only extract the first line of code generated by the model. Specifically, we allow the model to generate 64 tokens for each prompt and only take the content before the first newline character as the output. Then, we compare the generated code with both the buggy code and the fixed code, yielding three possible cases: (1) the output matches the buggy code, (2) the output matches the fixed code, or (3) it does not match either. Ideally, a model should generate more outputs that match the fixed code and fewer outputs that match the buggy code. By sending many prompts to a model, Jesse et al. count the number of cases (1) and (2) and compute their ratio to show that models tend to generate buggy code. As these models are non-deterministic, the same prompt may lead to different outputs. We let the model generate 10 outputs for each prompt and count the number of cases (1) and (2). In our experiment, we want to see how online and offline hotfixing can help models generate less buggy code and more fixed code.
We additionally consider hotfixing LLM4Code to mitigate the privacy information leakage issue as a case study. The process for constructing this dataset is explained in Section 5.
3.5. LLM4Code Selection and Training
@@
Considering the popularity and high performance of decoder-only models in code generators (Hou et al., 2024), we focus on decoder-only models in our evaluation. Following the study that exposes the undesired behaviors of code generators (Jesse et al., 2023), we choose CodeGen (Nijkamp et al., 2023) models as our experiment subject, which are open-source models and have been widely used when evaluating code generation models. The CodeGen architecture follows a standard transformer decoder with left-to-right causal masking. Nijkamp et al. (Nijkamp et al., 2023) train multiple CodeGen models using different configurations and datasets. Although CodeGen-mono achieves the best performance on the HumanEval benchmark, it only supports the generation of Python code. We choose CodeGen-multi, which can support multiple programming languages, as our experiment subject. We use three CodeGen-multi models of different sizes in terms of the number of parameters: 350M and 2B. [^5] In this paper, we use CodeGen-350M and CodeGen-2B to refer to the CodeGen-350M-multi and CodeGen-2B-multi. Considering that we need to train multiple models with different fine-tuning methods, we exclude the larger versions with 6 and 16 billion parameters due to the limits of our computational resources.
One key requirement for hotfixing LLM4Code is that the hotfixing should be time-critical, small in size, fast to obtain, and target a specific issue in LLM4Code. Thus, the conventional fine-tuning method (i.e., updating the entire model) is not suitable for hotfixing. We employ various parameter-efficient fine-tuning (PEFT) techniques to update the models, which can reduce the number of trainable parameters and thus reduce the computational cost. Following a recent study on evaluating PEFT on LLM4Code (Weyssow et al., 2024), we employ four PEFT techniques: LoRA (Hu et al., 2022), IA3 (Liu et al., 2022), Prefix tuning (Li and Liang, 2021), and QLoRA (Dettmers et al., 2023). For QLoRA, we quantize the model to 8-bit floating-point and 4-bit floating-point data types. In the following part of the paper, we use the acronyms FT, LoRA, IA3, Prefix, QLoRA-8bit, and QLoRA-4bit to represent the traditional fine-tuning, LoRA, IA3, Prefix tuning, QLoRA with 8-bit quantization, and QLoRA with 4-bit quantization, respectively. Note that the hotfix is in different forms depending on the PEFT methods used. For example, the hotfix using prefix tuning is a set of virtual tokens added to the input tokens of the LLM, and the hotfix using LoRA is a set of low-rank matrices added to the attention layers of the LLM.
We downloaded the CodeGen models from Hugging Face and ran them on two NVIDIA GeForce A6000 GPUs with 48 GB of memory. We extend the open-source repository provided by Weyssow et al. (Weyssow et al., 2024) to implement different fine-tuning methods with our proposed loss functions. Following their settings, we set the learning rate to 3e-4. For prefix-tuning, the prefix is set to be 20 trainable continuous tokens. Following recent studies (Weyssow et al., 2024; Li et al., 2024b), we trained CodeGen-350M for five epochs and CodeGen-2B for three epochs using different PEFT methods. The Adafactor optimizer (Shazeer and Stern, 2018) with 16-bit floating-point precision is used for training.
@@ 3.6.2. Perplexity
$N$ is the total number of words in the sequence. $P(w_{i}\mid w_{1},\ldots,w_{i-1})$ is the probability assigned by the model to the word $w_{i}$ given the preceding words $w_{1},\ldots,w_{i-1}$. A lower perplexity score indicates that the model’s knowledge is more consistent with the data used to compute the perplexity. We aim to evaluate how hotfixing may affect the model’s knowledge about the original training data. However, the training data of both CodeGen-350M and CodeGen-2B are not publicly available. Given the insights from previous studies that larger models can memorize much training data (yang2023memorization; Carlini et al., 2021), we sample 10,000 outputs from the CodeGen-16B model, which are expected to be similar to the training data of CodeGen-350M and CodeGen-2B. Thus, these sampled outputs are appropriate to evaluate how the model’s knowledge about the original training data is affected by hotfixing. We measure the average perplexity of CodeGen-350M, CodeGen-2B, and their hotfixed versions on the 10,000 outputs sampled from CodeGen-16B.
4. Results
This section presents three research questions to understand the feasibility of offline and online hotfixing to mitigate undesirable behaviors in LLM4Code.
RQ1. How do offline and online hotfixing perform?
RQ2. How do learning objectives affect offline hotfixing results?
RQ3. What side effects are potentially caused by hotfixing and how to mitigate the undesired side effects?
The first question compares offline and online hotfixing. It also tries different PEFT strategies while keeping the same learning objectives. The second question uses the same fine-tuning method (LoRA) and tries different learning targets to understand how they affect the hotfixing results. We also examine how hotfixing causes side effects and how to mitigate them.
Table 1. This table shows how PEFT methods affect hotfixing. We set the loss function of each model to be the same. The integer number in each cell represents the number of generated buggy/fixed code. The percentage in the parentheses shows the change compared to the original models’ results. $\downarrow$’ and $\uparrow$’ indicate the decrease and increase, respectively. The cells with the best results are highlighted in green.
@@ 4. Results
4.1. RQ1. How do offline and online hotfixing perform?
@@
In this research question, we evaluate both online and offline hotfixing. For offline hotfixing, we aim to understand which PEFT methods can perform better in hotfixing LLM4Code. Here, performance has two aspects: (1) effectiveness: to what extent the hotfix can make the model generate less buggy but more fixed code; and (2) efficiency: how many computational resources are required to hotfix the model.
@@
Table 1 shows the results of models before hotfixing and models hotfixed with online and offline methods. First, the third column shows the number of buggy code and fixed code generated by the original CodeGen-350M and CodeGen-2B models. We observe that both models tend to generate more buggy code than fixed code, confirming the findings in Jesse et al.’s study (Jesse et al., 2023). We also observe that the larger model generates more buggy code and fixed code than the smaller model, indicating that larger models have a stronger capacity to memorize the training data (Yang et al., 2024c). Specifically, CodeGen-2B generates 13.26% more buggy code and 39.69% more fixed code than CodeGen-350M. Then, we choose the mitigation method used in (Jesse et al., 2023) as an example of online hotfixing and find that adding comments to the model prompt can mitigate the issue to some extent, but the mitigation performance is limited. Specifically, this method only reduces 7.58% and 10.79% of buggy code, and increases 13.12% and 6.38% of fixed code for CodeGen-350M and CodeGen-2B, respectively. Note that this method does not incur additional cost for training the model.
We then evaluate the performance of online hotfixing using different PEFT methods. In this RQ, we use the loss function defined in Equation (5) to train the models (i.e., learning the added code and penalizing the removed code). Although using the same loss function, we notice that different PEFT methods can lead to different hotfixing results. In Table 1, we label a cell as green if a PEFT method achieves the best performance. We can observe that LoRA has 3 green cells, and the other one is QLoRA-4bit. It indicates that LoRA-based methods are more effective in hotfixing LLM4Code. Specifically, using LoRA as a hotfix can reduce 44.32% (for CodeGen-350M) and 50.47% (for CodeGen-2B) of buggy code, and increase 108.43% (for CodeGen-350M) and 103.38% (for CodeGen-2B) of fixed code. In contrast, IA3 and Prefix-tuning are less effective than LoRA. IA3 and Prefix-tuning are close in terms of reducing the generated bugs. For example, they reduce 36.80% and 36.45% of buggy code generation on CodeGen-2B, which is still lower than LoRA (48.35%). The gap between LoRA-based methods and the other two methods in guiding models to generate more fixed code is larger. For example, Prefix-tuning only leads to a 5.40% increase in fixed code generation for CodeGen-2B, and IA3 leads to 72.62%, which is less than the 103.38% increase brought by hotfixing with LoRA.
We also analyze the efficiency of different PEFT methods for hotfixing LLM4Code. We set the batch size to 1 for each method and record the average time required to conduct one epoch of learning on the same A6000 GPU. We repeat the experiment 5 times (setting the training epoch number as 3) and report the average time. On CodeGen-350M, LoRA and Prefix-tuning show better time efficiency, taking 16.8 and 16.2 minutes, respectively. IA3 takes 19.8 minutes, faster than QLoRA-4bit and QLoRA-8bit, which take 21.0 and 28.8 minutes, respectively. However, full parameter fine-tuning CodeGen-350M takes 71.1 minutes on the same dataset. Additionally, hotfixing using LoRA produces a set of new parameters that is relatively small in size (5.0 MB for CodeGen-350M and 13.9 MB for CodeGen-2B). This means that applying the hotfix does not require transferring the entire model, which can be significantly larger in size (5.69 GB for a CodeGen-2B). This suggests that hotfixing using LoRA is fast to obtain and small in size.
Answers to RQ1: Offline hotfixing using parameter-efficient fine-tuning (PEFT) demonstrates better results than online hotfixing using prompt engineering. Hotfixing results vary for different PEFT methods. LoRA-based methods demonstrate stronger effectiveness than IA3 and prefix-tuning. In our best setting, hotfixing can reduce 44.32% and 50.47% of buggy code and increase 108.43% and 103.38% of fixed code for CodeGen-350M and CodeGen-2B.
4.2. RQ2. How do learning objectives affect offline hotfixing results?
In RQ1, we find that offline hotfixing with LoRA demonstrates stronger effectiveness than other PEFT methods. In RQ2, we use LoRA to train models and analyze how different learning objectives affect hotfixing results. In Section 3.3, we define four settings of loss functions, named Vanilla, Guided, Unlearn, and Dual loss settings. In the Vanilla setting, we minimize the loss on the entire example after the fix, i.e., the commonly used setting in fine-tuning LLM4Code (Weyssow et al., 2024), defined in Equation (2). In the Guided setting, we guide the model to minimize the loss only on the added fixed code, defined in Equation (3). In the Unlearn setting, we guide the model to maximize the loss on the deleted code, defined in Equation (4). In the Dual setting, we additionally penalize the loss on the deleted code on top of the Guided setting, defined in Equation (5).
@@
Table 2. The table shows how loss functions affect hotfixing. We use LoRA to train each model. The percentage number under ‘Vanilla’, ‘Guided’, ‘Unlearn’, and ‘Dual’ columns are changes compared to the original model. $\downarrow$’ and $\uparrow$’ indicate the decrease and increase.
Type Models Vanilla Guided Unlearn Dual
@@
The results of each loss setting are shown in Table 2. We can observe that the Unlearn setting makes the model lose its ability to generate meaningful code (the hotfixed models complete zero buggy code and zero fixed code), which is not desired in hotfixing. It suggests that directly maximizing the loss on the undesired code is not effective in hotfixing. It can be observed that the other three settings (Vanilla, Guided, and Dual) can mitigate the undesired behaviors of LLM4Code. The Vanilla setting reduces 11.38% and 8.69% of buggy code and increases 64.89% and 48.02% of fixed code for CodeGen-350M and CodeGen-2B, respectively. By guiding the model to specifically learn from the added fixed code, the Guided setting can reduce 37.81% and 41.88% of buggy code and increase 93.11% and 100.90% of fixed code for the two models. Additionally penalizing the model on the deleted code in the Dual setting can further improve the performance. Taking CodeGen-350M as an example, the Dual setting can reduce 41.18% of buggy code and increase 107.05% of fixed code, outperforming the Guided setting that reduces 37.81% of buggy code and increases 93.11% of fixed code. The results suggest that guiding the model to learn the difference between two versions of the dataset is more effective in hotfixing.
Table 3. This table shows how different loss function settings with KL divergence affect the results of hotfixing. We set the fine-tuning method of each model to be the same, i.e., LoRA. The results are the average of five runs. The percentage numbers under the ‘Vanilla+KL’, ‘Guided+KL’, and ‘Dual+KL’ columns are the changes compared to their counterparts without KL divergence.
Type Models Loss Functions Average Changes Caused by KL Loss
Vanilla+KL Guided+KL Dual+KL
No. of SStuBs 350M
2,687
(3.03% $\uparrow$)
@@
We also investigate the impact of adding KL divergence to the loss function. The goal of adding KL divergence is to reduce the potential side effects caused by hotfixing (detailed analysis in RQ3). However, this may also affect the hotfixing results, and we show that such impacts on hotfixing results are limited. Table 3 shows the results when adding KL divergence to each loss function, in columns with the suffix ‘+KL.’ The percentage number in this table is the change compared to the results without KL divergence. We find that adding KL loss does not have a significant impact on the hotfixing results. We report the average changes caused by KL loss in the last column of Table 2. On average, adding KL loss only increases 3.41% of buggy code and reduces 1.51% of fixed code for CodeGen-350M, and increases 2.74% of buggy code and reduces 1.57% of fixed code for CodeGen-2B, compared to the results without KL loss. However, as we will show in the next research question, adding KL loss can help mitigate the undesired side effects caused by hotfixing LLM4Code.
@@
Answers to RQ2: The Dual loss setting can achieve the best performance, reducing 41.18% and 50.47% of buggy code and increasing 107.05% and 103.38% of fixed code for CodeGen-350M and CodeGen-2B, respectively. Adding KL divergence to the loss function does not have a significantly negative impact on results.
@@
4.3. RQ3: What side effects are potentially caused by hotfixing and how to mitigate the undesired side effects?